Page tree
Skip to end of metadata
Go to start of metadata

Lab-as-a-Service (LaaS) Meetings

Logistics

  • Weekly meetings:
    • Weekly: Tuesday 9-9:30am PT / 6-6:30pm CET (with audio)
  • To join the meeting:
    • GTM/Webex: details to be shared over IRC at the start of the meeting.


 

January 16, 2018

Agenda:

  • Update on ARM servers (Ray, Lincoln)
  • Lab status (Lincoln)
  • MVP status (Lincoln)

Minutes:

  • ARM servers shipment: No immediate update. Ray will follow up with the vendor to see whether shipment is about to happen.
  • 40G cables: Should have shipped today
  • 14 X710 NICs were returned to CDW.
  • Lab status:
    • Racks still haven't arrived (they are supposed to arrive today)
  • MVP status:
    • System is ready to work with the HP servers, once they are racked we should be ready to roll.
    • Plan to do a demo next week.
  • ONS
    • Fathi and Jack submitted a talk about LFN Infra and LaaS: LaaS should see some promo in several sessions at ONS.

Recording: https://cisco.webex.com/ciscosales/lsr.php?RCID=0356930cd048471091d2c7a9138bc478 (pw: pEkFSje8 )

January 9, 2018

Agenda:

  • Update on ARM servers (Ray, Lincoln)
  • Use of spare X710 NICs
  • Lab status (Lincoln)
  • MVP status (Lincoln)

Minutes:

  • ARM server shipment: Ray gave an update via email:
    • The following is the delivery schedule after consulting with with folks like Joe/Bob/Alexandru...  T60* are T61 models with reconfigured firmware

      Convert T61 to T60*
      US Stock
      5 x R270-T60: 1/17/18
      4 x R270-T60*           1/31/18
      GBT Taiwan Stock
      5 x R270-T60*           2/16/18

  • Use of spare X710 NICs
    • We'll return 14 NICs to CDW (only 14 can we returned, because we'll need to at least purchase 100 to keep the price).
    • We'll consider deploying some ARM servers with removed cover, so that the full-height NICs can be slotted in.
      • This approach will be given a test-drive once ARM servers are available. Air-flow etc. might be an issue.
  • Lab status
    • Rack delivery was delayed due to weather. Racks are assumed to arrive tomorrow (Jan/10). 
      • This causes ~ 1 week delay of our schedule towards MVP (new ETA for MVP: Jan/26)
    • 40G cables were also ordered and are expected to arrive this week as well
  • MVP/Dashboard
    • All Laas dashboard JIRA tickets also have a "dashboard" label now (thanks to Jack).
      This label can be used in addition to the "Laas" component to provide for filtering etc.
    • Parker is out this week - hence limited progress. He cleaned up a lot of issues seen already. Steady progress has been made.

 

January 2, 2018

Agenda:

  • Summary of equipment procurement/order status (Lincoln, Ray)
  • Review updated lab cabling plan (Lincoln)
  • Review MVP tasks (are there tasks which don't have an owner yet)?

 

Minutes:

  • Equipment
    • All servers and switches (except the ARM servers) have arrived at UNH
    • Racks become available on Thursday this week. Equipment will be racked and cabled from Thursday onwards.
    • 40G interconnect cables are about to be ordered
    • No new switch required (different from what we thought just before the break: The ARM servers only have 4 ports. The Intel NICs are full height and won't fit the 1/2 height ARM server form factor. This frees up some ports, hence no additional switch is needed)
  • Cabling plan

 

December 26, 2017

No meeting (holiday break)

December 19, 2017

Agenda:

  • Summary of equipment procurement/order status (Ray)
  • Close on open questions for Lab cabling plan (Lincoln)
  • Review JIRA tickets for MVP (Portal etc.) - EPIC:  PHAROS-318 - Getting issue details... STATUS
    • Assign/decide owners for tickets

 

Minutes: http://ircbot.wl.linuxfoundation.org/meetings/opnfv-meeting/2017/opnfv-meeting.2017-12-19-17.00.html

December 12, 2017

Agenda:

  • Summary of equipment procurement/order status (Ray)
  • Review status of cabling (Lincoln)
  • Review JIRA tickets for MVP (Portal etc.) - EPIC:  PHAROS-318 - Getting issue details... STATUS
    • Assign/decide owners for tickets

Minutes: 

  • Summary of equipment status: All Cisco gears shipped.  Intel NICs are on back order, and Ray to follow-up with CDW.  Lincoln added that Racks and power strips have been ordered by UNH.  After the meeting, received an email that HPE equipments are shipping and Ray will get the tracking info.  
  • Review status of cabling: Lincoln uploaded a couple of documents at  PHAROS-335 - Getting issue details... STATUS  that include cabling diagram.  Consensus was to continue the discussion via the JIRA ticket.
  • Review of JIRA tickets for MVP:  Didn't have nearly enough time to discuss the topic, but people noted that we need to add new capabilities to the existing LaaS dashboard.  A suggestion was made to have an hour long meeting next week to continue discussion on both cabling and open items for the MVP
  • "Governance" of the LaaS initiative:  Consensus to keep this initiative within the Pharos project.  

 

December 5, 2017

Agenda:

  • F2F LaaS team meeting at Plugfest (date/time/room) (Ray)
  • Summary of equipment procurement/order status (Ray)
  • Review status of cabling (Lincoln)
  • Review & prioritize JIRA tickets for MVP (Fatih)
  • Installer integration (Parker)

Minutes: http://ircbot.wl.linuxfoundation.org/meetings/opnfv-meeting/2017/opnfv-meeting.2017-12-05-17.05.html

November 28, 2017

Agenda: 

  • Decide on weekly meeting time
  • Summary of equipment procurement/order status (Ray)
  • Summary of the UNH lab status & projected timelines for equipment availability, once the equipment has been delivered to UNH (Lincoln)
  • Brief overview/summary of current LaaS software status (portal etc.)
  • Define the first offering of LaaS (aka “minimum viable product”) that LaaS would start off with
  • Define next steps & owners

Recording: https://cisco.webex.com/ciscosales/lsr.php?RCID=d402e776e8574d07acb8da4d2ed28417 (pw: TwkEM9pY)

Brief minutes:

  • Weekly meetings: Tuesday, 9-9:30 am PT
  • Equipment: Expected arrival at UNH is around Christmas. UNH will have someone on site to receive the equipment.
  • Rack & cable up: Expected to happen 1st week of January (if equipment is received by the end of 2017)
  • UNH already has pieces of the solution available – e.g. bare metal automated install (for Ubuntu, Centos); OPNFV virtual deploys (Fuel, Joid).
    Portal is lagging behind - https://wiki.opnfv.org/display/INF/Dashboard+Roadmap – Portal was built as an intern project and requires a set of enhancements to be useable for LaaS.
    Portal is hosted by LF right now. Now major benefits to change this. We’ll keep things as they are for now.
  • Initial offering / MVP:
    • Automated install of a single bare metal server. User should be able to choose from at least 3 images (covering the main distros: Centos, Ubuntu, Suse)
    • Reservation will follow fixed timeslots (e.g. 1 week). All reservations will initially be synched to allow for easier system maintenance (e.g. all reservations end always on Sunday, midnight PT).
    • Next steps:
      • AI Lincoln: Create cabling plan for server/switches
      • AI Fatih: Create Jira tickets for tasks (mostly focused on Portal) – using https://wiki.opnfv.org/display/INF/Dashboard+Roadmap as starting point.
        We plan to use Jira components. Tickets will be created under Pharos project but should differentiate LaaS and Dashboard.
      • AI Ray: Find a room/time for F2F discussions at the hackfest.
      • Next meeting: Tuesday Dec/5, 9am PT – Frank to send out invite/webex.
  • No labels